OpenStudy (dezandalf42714):

Ok so I tried this question on my own and got 31 1/4 but it says im not right any help pls? I tried everything but nothing is working. I know you are supposed to work backwards but I cant find an answer. Jake's band played songs at three different parties. At the first party they played one-fifth as long as they did at the second party, which was one and one-fourth as long as the third party. If they played for 5 minutes at the first party, how many minutes did they play at the third party? Select one: A. twenty minutes B. twenty-six and one-fourth minutes C. twenty-eight minutes

OpenStudy (anonymous):

you still there?

OpenStudy (dezandalf42714):

yes

OpenStudy (anonymous):

k cool the answer is c

OpenStudy (dezandalf42714):

ok. can you explain how you got that?

OpenStudy (welshfella):

Time at first party = 1/5 x (x = time at "2nd party) 1/5 x = 5 so x = 5 / (1/5) = 25 minutes

OpenStudy (dezandalf42714):

ok thats what I got for the second party. 25 minutes

OpenStudy (welshfella):

Now this was 5/4 times as long as the 3rd party So time in 3rd party = 25 / (5/4) = 25 * 4/5 = 20 minutes

OpenStudy (dezandalf42714):

Oooooh ok!! Thanks a bunch!!!

OpenStudy (welshfella):

Yw